What are the solutions of the equation (2x - 7)( x + 10) = 0 ?
A -3.5 and 10
B 7 and –10
C 2/7 and –10
D 3.5 and -10

Answer by ewatrrr(24785)

Hi.
.
(2x - 7)( x + 10) = 0
.
(2x - 7)= 0
x = (7/2) or 3 1/2 or 3.5
.
( x + 10)= 0
x = -10
D would be the correct selection